mr. whitney is three times as old as his son. twelve years from now he will only be twice as old. what are their ages now?
Answer by jorel1380(3719)

W=3S
W+12=2(S+12)
3S+12=2S+24
S=12
W=36
Mr. Whitney is 36. His son is 12.